OpenStudy (anonymous):

integrate[sqrtx/(x-4)

terenzreignz (terenzreignz):

\[\Large \int \frac{\sqrt x}{x-4}dx\]

OpenStudy (anonymous):

use substitistion man

terenzreignz (terenzreignz):

He has already left...

OpenStudy (anonymous):

i knw i just doing when he comes may take his answer

OpenStudy (anonymous):

u will have (X^1/2)^2-2^2 and this is how u r goona substitue x^1/2=t and then t/t^2-2^2 nw using substitution u can solve it

OpenStudy (anonymous):

also u hv to change dx to dt remember it ok

OpenStudy (anonymous):

use x=t^2 dx =2tdt the given question becomes 2t^2/(t^2-4)=2*(t^2-4 +4)/(t^2-4)=2* (1 + 4/(t^2-4)) which can be integrated easily